# Activation Steering for Robust Type Prediction in CodeLLMs

### v1

@@ Abstract
Contemporary LLMs pretrained on code are capable of succeeding at a wide variety of programming tasks. However, their performance is very sensitive to syntactic features, such as the names of variables and types, the structure of code, and presence of type hints. We contribute an inference-time technique to make CodeLLMs more robust to syntactic distractors that are semantically irrelevant. Our methodology relies on activation steering, which involves editing internal model activations to steer the model towards the correct prediction. We contribute a novel way to construct steering vectors by taking inspiration from mutation testing, which constructs minimal semantics-breaking code edits. In contrast, we construct steering vectors from semantics-preserving code edits. We apply our approach to the task of type prediction for the gradually typed languages Python and TypeScript. This approach corrects up to 90% of type mispredictions. Finally, we show that steering vectors calculated from Python activations reliably correct type mispredictions in TypeScript, and vice versa. This result suggests that LLMs may be learning to transfer knowledge of types across programming languages.
@@ 1 Introduction
Large Language Models trained on code (CodeLLMs) are one of the most successful applications of LLMs to date. They are the foundation for both developer tools (e.g., Weiss & Yahav (2013); git (2021)) and for reasoning agents (e.g., Zelikman et al. (2024); Yang et al. (2024); Hu et al. (2024); Li et al. (2022)). However, despite their utility, CodeLLMs can still be unreliable. Recent work shows that even the most capable models are not robust to small semantics-preserving changes such as refactorings and variable renamings (Hooda et al., 2024; Tambon et al., 2024). Thus CodeLLMs often cannot be deployed without careful supervision, such as a human-in-the-loop or robust code sandboxing.
In this paper, we present an approach to make CodeLLMs more robust using activation steering (Li et al., 2024; Rimsky et al., 2023; Turner et al., 2023; Subramani et al., 2022). Activation steering is an inference-time model editing technique that modifies the intermediate computations (activations) of a model using steering vectors to steer model behavior toward desired outcomes. We present a novel technique for constructing steering vectors for the code domain by taking inspiration from mutation testing (DeMillo et al., 1978). Mutation testing automatically constructs minimal semantics-breaking code edits. In contrast, we construct steering vectors from minimal semantics-preserving code edits that lead to CodeLLM mispredictions. The nature of code allows us to construct these code edits in a sound and scaleable way.
@@
[…]
Finally, we are surprised to find that type prediction steering transfers across programming languages. In particular, the steering vectors that we construct from TypeScript prompts are very effective at steering Python type predictions, and vice versa. We conduct a careful analysis and conclude that is it likely that multi-lingual CodeLLMs are learning a representation of types that is shared across the programming languages we study (§ 4.5).
2 Background and Related Work
@@
CodeLLMs are capable of performing a variety of downstream tasks such as generating code from natural language, explaining code, generating tests, and more (Nam et al., 2024; Schäfer et al., 2024). Contemporary CodeLLMs are decoder-only transformers trained on vast amounts of source code (Li et al., 2023; Rozière et al., 2024; Guo et al., 2024). The aforementioned models are also trained to fill-in-the-middle (FIM) (Bavarian et al., 2022; Fried et al., 2023). FIM training does the following (1) It splits $\approx 50\%$ of training items into three chunks—prefix, middle, and suffix—of random lengths; (2) It adds special tokens to the start of each chunk to demarcate chunk boundaries; and (3) It reorders the middle chunk to appear last. The language modelling training objective remains unchanged. Thus FIM inference allows us to generate the middle chunk, conditioned on the prefix and the suffix (figure 1). The resulting model can still be used for conventional left-to-right generation.
@@
In this paper we study activation steering for type prediction, formulated as a fill-in-the-middle task. Prior work uses this formulation to evaluate base models (Yee & Guha, 2023; Fried et al., 2023), but does not study models’ internal mechanisms or try to improve task performance in a targeted way. There is prior work that trains smaller, specialized models for type prediction (Hellendoorn et al., 2018; Jesse et al., 2022; 2021; Pandi et al., 2021; Wei et al., 2020), but contemporary CodeLLMs outperform these specialized models (Yee & Guha, 2023; Fried et al., 2023).
Classical type prediction and type inference
Type prediction is also known as type migration, and is distinct from type inference as found in languages such as OCaml and Haskell. In those languages, every variable is typed, even if the types are implicit (Harper & Mitchell, 1993). In contrast, gradual type prediction can change program semantics (Phipps-Costin et al., 2021). There is work on gradual type prediction that uses human-written constraints and constraint solvers (Phipps-Costin et al., 2021; Rastogi et al., 2012; Siek & Vachharajani, 2008; Campora et al., 2018; Henglein & Rehof, 1995; Cartwright & Fagan, 1991). But, these papers present algorithms for variations of the lambda calculus or simple functional languages such as Scheme, and have not been scaled to more complex, modern programming languages.
@@
Interpretability seeks to understand the causal mechanisms behind model inference. In particular, mechanistic interpretability studies how model behaviors arise from specific components within the transformer architecture. One of the driving motivations of the field is that by understanding how a model works, we can understand how to align models to desired behavior.
Prior work in mechanistic interpretability includes localizing and editing factual associations within the transformer (Meng et al., 2022), as well as probing hidden representations for knowledge of high-level concepts (Li et al., 2024; Dong et al., 2023). These works are forms of implicit evaluation of model ability, as opposed to explicit evaluation using benchmarks (Dong et al., 2023). Mechanistic interpretability has contributed a wealth of methods for implicit evaluation, for example activation patching (Vig et al., 2020; Variengien & Winsor, 2023), which applies patches to model activations in order to produce a change in behavior. Research leveraging activation patching has suggested the existence of task vectors (Hendel et al., 2023; Ilharco et al., 2022). Task vectors are representations within the transformer which encode a high-level notion of the task described by the prompt. The existence of task vectors is further supported by work in activation steering. Steering has been employed to diminish model deceitfulness and sycophancy (Rimsky et al., 2023; Li et al., 2024), which suggests that steering works by modifying the activations corresponding to task vectors. We thus draw on the idea of task vectors to perform our steering experiments. Importantly, steering relies on the linear representation hypothesis, which argues that concepts are represented as directions in the embedding space of a model (Park et al., 2023). We posit that successful steering performs transformations over model activations that shift the task representation towards a target direction.
Despite much work on the mechanistic interpretability of LLMs, similar research on CodeLLMs has been limited. In this paper we focus on types, which are a fundamental feature of programming languages. One obstacle in this domain is that within a program, a type annotation is semantically constrained by both the tokens that precede it and those that succeed it. For example, the type of a function argument is constrained by how the argument is subsequently used in the function body. To overcome this, we use CodeLLMs that are trained on FIM. This allows us to construct prompts for type prediction.
Mutation testing and program transformations
We take inspiration from mutation testing (DeMillo et al., 1978), but our technique is different. The goal of mutation testing is to test a program’s test suite. To do so, a mutator injects small bugs that alter the semantics of a program, such as changing a 00 to a $1$ or turning $x>y$ into $x<y$. The hypothesis is that a good test suite should be able to catch these artificial bugs, and there is a substantial evidence that the ability to catch both artificial and real-world bugs is strongly correlated (Just et al., 2014). In this paper, we make program edits that would not affect test cases. [^1] Reflection in TypeScript and Python allows a program to detect any source code change. But, most test suites do not use these features. Instead, we make syntactic edits that lead to type mispredictions for a given CodeLLM.
@@ 3.1 Constructing Steering Datasets
We build steering datasets for 1B and 7B parameter StarCoderBase models. These models are trained to fill-in-the-middle (§ 2), which is necessary for the type prediction task.
@@
A distinguishing feature of our work is that we construct steering pairs from natural data (source code from GitHub) instead of using templates to construct simple examples. This provides more evidence that activation steering is robust. For Python, we use ManyTypes4Py (Mir et al., 2021), a dataset for neural type prediction. It features code from 5,382 Python projects that use Python type annotations and successfully type-check. For TypeScript, we filter The Stack (Kocetkov et al., 2023) to find 1.1M TypeScript files that type-check, of which we sample up to 2,000 for steering.
@@
Figure 3: A fragment of a Python steering pair. The original code is 70 lines of text. The dict is the expected prediction. But, renaming config to __tmp0 makes the model mispredict Repository, which is a hallucinated type.
@@
For a given model $M$, our steering dataset is a set of triples $(x^{+},x^{-},t)$ where $t$ is a type and $x^{+},x^{-}$ are prompts that represent fill-in-the-middle type prediction tasks with $t$ as the expected type for both prompts. However, the maximum likelihood generation for $x^{+}$ is $M(x^{+})=t$ and for $x^{-}$ it is $M(x^{-})=k$, where $k\neq t$.
@@
To construct the negative prompt, we incrementally apply semantics-preserving code edits to the positive prompt until $M$ mispredicts $t$. We implement the following edits. 1) Rename variable:We select a variable binding (from a function definition) and rename it to an arbitrary name that does not conflict with other variables in the program. We also rename all bound occurrences of the same variable so that the program’s semantics remain unchanged. 2) Remove type annotation:We select a type annotation (excluding the target $t$) and delete it. In a gradually typed language, removing or relaxing an annotation does not alter program semantics. 3) Rename user-defined type:We select an arbitrary type definition (e.g., a class name or a type alias) and rename it to an arbitrary name that does not conflict with other names in the program. 4) Rename builtin type:We also support renaming builtin types by introducing type aliases. Figure 2 illustrates several separate edits to a program.
By incrementally applying edits, the model eventually mispredicts $t$ (or else we fail to find an $x^{-}$ and thus discard $x^{+}$). Figure 3 illustrates a real example from our dataset where StarCoderBase-1B mispredicts after an edit. Note that a single edit often alters the positive prompt $x^{+}$ at several points. Moreover, when we apply several edits, the two prompts $x^{+},x^{-}$ may become significantly different from each other. We find that the mean edit-distance for our Python and TypeScript steering pairs is $74$ and $327$ respectively. This is in contrast to steering pairs constructed from sentence templates, that only differ by a small, fixed number of words at a far smaller number of places.
We hypothesize that mispredictions happen because the model fails to identify the inference task as a type prediction task. This likely occurs because our edits shift the prompt outside the distribution of a model’s training data. We thus call our edits distractors, since they distract from the correct task. We leverage activation steering to make the model more robust against such distractors and other semantically irrelevant features of code. For example, distractors often remove important in-context clues like function names which models rely on for prediction. Relying on textual meaning rather than program structure is a limitation of CodeLLMs that we improve with activation steering.
Class balance
@@
Table 1: Overview of StarCoderBase-1B dataset sizes and applied code edits. Note that steering subsets are used for constructing steering vectors, while held out subsets are used to evaluate steering performance on out of distribution data. Dataset sizes vary because we only apply edits to programs where the model correctly predicts types. We then filter to remove overlap in source programs between held-out and steering sets. This accounts for the variation in sizes.
@@
We construct several datasets $\mathcal{D}$, where each dataset applies a subset (or all) of the program edits described above to programs in a language. We further split each dataset into a steering subset and a held-out evaluation subset. We ensure that the two subsets have prompts based on different source programs, to test how steering generalizes to unseen programs. Table 1 summarizes the dataset sizes for StarCoderBase-1B. We report the sizes for StarCoderBase-7B in appendix A.
@@ 3.2 Constructing Steering Vectors
Given a dataset of steering pairs $(x_{i}^{+},x_{i}^{-},t)\in\mathcal{D}$ and model $M$, we apply a forward pass to every $M(x_{i}^{+}),M(x_{i}^{-})$ to collect model activations on the last token. Specifically, we extract activations from the residual stream of the model, which is the output of a transformer layer $\ell_{i}$ that in turn becomes the input to subsequent layers $\ell_{i+1...n}$. We write $A_{\ell}(x)$ to denote this activation tensor at some layer $\ell$ for prompt $x$. Thus we compute steering vectors $\mathbf{t}_{\ell}$—one for each layer—as the mean difference between positive and negative activations at that layer:
@@
The intuition behind equation 1 is that the distance between positive and negative pairs in activation space encodes the transformation for steering towards the correct type.
@@
4 Evaluation
4.1 Layer Ablations

### v2

# Understanding How CodeLLMs (Mis)Predict Types with Activation Steering

@@ Abstract
CodeLLMs are transforming software development as we know it. This is especially true for tasks where rule-based approaches fall short, like type prediction. The type prediction task consists in adding a new type annotation to a partially typed program, such that the resulting program is closer to being fully typed. The intractability of rule-based approaches and high cost of manual annotation make CodeLLMs an attractive solution to the problem. However, CodeLLMs are still far from being deployed on the large-scale due to doubts surrounding their reliability.
To shed some light on how CodeLLMs approach type prediction, we investigate what happens when a model mispredicts a type. We show that by applying semantics-preserving edits to code, CodeLLMs are eventually misled into mispredicting type annotations. However, by leveraging activation steering we are able to “steer” the model back to the correct prediction, making models more robust against semantically irrelevant prompt features. We show that steering achieves comparable performance to fine-tuning directly on the type prediction task. Furthermore, we find that steering vectors computed from Python code are effective at correcting TypeScript mispredictions, and vice versa. To our knowledge, this is the first evidence of its kind to suggest that CodeLLMs learn task representations that transfer across languages.
Understanding How CodeLLMs (Mis)Predict Types with Activation Steering
Francesca Lucchetti & Arjun Guha Khoury College of Computer Sciences Northeastern University Boston, MA 02115 {lucchetti.f,a.guha}@northeastern.edu
@@ 1 Introduction
Large language models trained on code, or CodeLLMs, are one of the most successful applications of LLMs to date. They are the foundation of both developer tools (e.g., Weiss and Yahav (2013); GitHub (2021)) and reasoning agents (e.g., Zelikman et al. (2024); Yang et al. (2024); Hu et al. (2024); Li et al. (2022)), capable of performing a variety of programming tasks across several programming languages. This paper focuses on type prediction for gradually typed programming languages, specifically Python and TypeScript. Given a partially typed program $p$ written in a gradually typed language, the type prediction task is defined as follows (Migeed and Palsberg, 2020):
@@
A gradually typed programming language allows programs to freely mix typed and untyped code, giving programmers more flexibility than traditional static typing affords (Siek and Taha, 2006; Tobin-Hochstadt and Felleisen, 2006). However, this flexibility comes with a downside: at runtime, untyped code still needs to type-correct for the program to run correctly. With omitted or weak type annotations, type errors may not be caught until program execution.
Automating type prediction is important for both scientific and practical reasons. From a practical perspective, well-known companies like Airbnb, Dropbox, Slack, Netflix, and many others have each taken several years to manually add type annotations to their multi-million line gradually typed codebases (Rudenko, 2020; Lehtosalo, 2019; Felix Rieseberg, 2017; Luke Autry, ; Sumana Mohan et al., 2022; Abacus, 2019; Mihai Parparita, 2020; Jake Zimmerman, 2022). Automated gradual type prediction would save significant effort and resources.
From a scientific perspective, type prediction is intractable, and rule-based approaches have not been scaled up to contemporary languages (§ 2). Because of this, prior work has explored leveraging CodeLLMs for type prediction. However, CodeLLM performance on the task is far from being able to automate large-scale type annotation efforts. In particular, we show that CodeLLM type predictions are sensitive to small semantics-preserving changes in prompts like renaming variables, which limits their practical utility.
The sensitivity of CodeLLMs towards trivial prompt changes begs the question of whether CodeLLMs approach type prediction in the “correct” way–that is, by inferring the type of a variable based on the operations performed over it. Alternatively, CodeLLMs could be relying on text features to make spurious predictions–for example, if a variable name is n, then its type is integer. We term these features “distractors”. As CodeLLM assistants are increasingly integrated into our workflows, understanding whether their predictions are principled or spurious is fundamental for security and large-scale adoption of these systems.
To answer the question of how CodeLLMs approach type prediction, we design the following experimental setting.
Given a type prediction prompt that the CodeLLM predicts correctly, we apply edits until the model output fails type checking. We construct our edits such that they do not matter for type prediction. To do so, we take inspiration from mutation testing (DeMillo et al., 1978). We present a variation of mutation testing that constructs minimal, semantics-preserving edits that trigger CodeLLM mispredictions. The nature of code allows us to construct these edits in a sound and scaleable way (§ 3.1).
Since the model could originally predict the prompt, but fails with our trivial edits, this could mean that CodeLLM type prediction is spurious and based on distractors. However, our findings suggest a different story: that mispredictions happen because the model fails to identify the inference task as a type prediction task. This likely occurs because our edits shift the prompt outside the distribution of a model’s training data. We show that by editing intermediate computations (activations) of the model, we can recover performance on type prediction despite the presence of distractors.
We employ activation steering (Li et al., 2024; Rimsky et al., 2023; Turner et al., 2023; Subramani et al., 2022) to steer model activations toward the target task of type prediction. Activation steering is an inference-time model editing technique that modifies the activations of a model using steering vectors to steer model behavior toward desired outcomes.
Finally, we show that steering transfers across programming languages: the steering vectors we construct from TypeScript code are very effective at steering Python type predictions, and vice versa. This suggests that there exists a robust, generalizable mechanism for type prediction in CodeLLMs which, when activated, can correctly predict types despite the presence of distractors. From this, we conclude that CodeLLMs are learning task-solving mechanisms that transfer across languages (§ 4.4). To our knowledge, this is the first empirical evidence of its kind that identifies the occurrence of language transfer within LLMs  [^1] Transfer knowledge has previously been observed in models that were trained/fine-tuned on a single language but perform well on cross-lingual tasks. However, training data contamination challenges the validity of this result..
@@
2 Background and Related Work
Classical type prediction and type inference
Type prediction is also known as type migration, and is distinct from type inference as found in languages such as OCaml and Haskell. In those languages, every variable is typed, even if the types are implicit (Harper and Mitchell, 1993). In contrast, gradual type prediction can change program semantics (Phipps-Costin et al., 2021). There is work on gradual type prediction that uses human-written constraints and constraint solvers (Phipps-Costin et al., 2021; Rastogi et al., 2012; Siek and Vachharajani, 2008; Campora et al., 2018; Henglein and Rehof, 1995; Cartwright and Fagan, 1991). But, these papers present algorithms for variations of the lambda calculus or simple functional languages such as Scheme, and have not been scaled to more complex, modern programming languages.
@@ 2 Background and Related Work
CodeLLMs are capable of performing a variety of downstream tasks such as generating code from natural language, explaining code, generating tests, and more (Nam et al., 2024; Schäfer et al., 2024). Contemporary CodeLLMs are decoder-only transformers trained on vast amounts of source code (Li et al., 2023; Rozière et al., 2024; Guo et al., 2024). The aforementioned models are also trained to fill-in-the-middle (FIM) (Bavarian et al., 2022; Fried et al., 2023). FIM training does the following (1) It splits $\approx 50\%$ of training items into three chunks—prefix, middle, and suffix—of random lengths; (2) It adds special tokens to the start of each chunk; and (3) It reorders the middle chunk to appear last. The language modelling training objective remains unchanged. Thus FIM inference allows us to generate the middle chunk, conditioned on the prefix and the suffix (figure 1). The resulting model can still be used for conventional left-to-right generation.
@@
We follow prior work in formulating type prediction as a fill-in-the-middle task  (Yee and Guha, 2023; Fried et al., 2023). Within a program, a type annotation is semantically constrained by both the tokens that precede it and those that succeed it. For example, the type of a function argument is constrained by how the argument is subsequently used in the function body. FIM allows us to construct prompts for type prediction as next token prediction, moving both function body and signature to the prompt prefix. Prior work also includes finetuning smaller, specialized models for type prediction (Hellendoorn et al., 2018; Jesse et al., 2022, 2021; Pandi et al., 2021; Wei et al., 2020), but contemporary CodeLLMs outperform them.
It is now well known that even the most capable CodeLLMs are sensitive to small variations in prompts (Hooda et al., 2024; Tambon et al., 2024). Prior work uses a black-box approach to study this phenomena. In contrast, we investigate type prediction in depth using a whitebox approach, to shed light on the causes of this sensitivity. We further investigate editing internal model components to improve models’ robustness against prompt variations.
Mutation testing and program transformations
To build our semantics-preserving edits, we take inspiration from mutation testing (DeMillo et al., 1978). The goal of mutation testing is to test a program’s test suite. To do so, a mutator injects small bugs that alter the semantics of a program, such as changing a 00 to a $1$ or turning $x>y$ into $x<y$. The hypothesis is that a good test suite should be able to catch these artificial bugs, and there is a substantial evidence that the ability to catch both artificial and real-world bugs is strongly correlated (Just et al., 2014). Our technique differs from mutation testing in one key way: we make program edits that would not affect test cases. [^2] Reflection in TypeScript and Python allows a program to detect any source code change. But, most test suites do not use these features. Instead, we make syntactic edits that lead to type mispredictions for a given CodeLLM.
@@
Mechanistic interpretability has contributed a wealth of methods for investigating models’ internal mechanisms, for example activation patching (Vig et al., 2020; Variengien and Winsor, 2023). Recently, research leveraging activation patching has suggested the existence of task vectors (Hendel et al., 2023; Ilharco et al., 2022)–latent representations within the transformer which encode a high-level notion of the task described in the prompt. The existence of task vectors is further supported by work in activation steering. Steering has been employed to control qualities like deceitfulness and sycophancy in model outputs (Rimsky et al., 2023; Li et al., 2024), which suggests that steering works by modifying the activations corresponding to task vectors. The connection between task vectors and steering is what motivates our investigation into steering for understanding type prediction.
Despite much work on the mechanistic interpretability of LLMs, similar research on CodeLLMs has been limited, and research using real-world (e.g. GitHub) code is even more so. This is because MI techniques are hard to adapt across different prompt formats Hobbhahn et al. (2022). We are able to overcome this by using FIM tokens to standardize our patching experiments.
@@ 3.1 Constructing Steering Datasets
We build steering datasets for 1B and 7B parameter StarCoderBase models Li et al. (2023). These models are both trained to fill-in-the-middle (§ 2).
@@
A distinguishing feature of our work is that we construct steering pairs from natural data (source code from GitHub) instead of using prompt templates to construct simple examples. This is to ensure that activation steering is robust across varied examples. For Python, we use ManyTypes4Py (Mir et al., 2021), a dataset for neural type prediction. It features code from 5,382 Python projects that use Python type annotations and successfully type-check. For TypeScript, we filter The Stack (Kocetkov et al., 2023) to find 1.1M TypeScript files that type-check, of which we sample up to 2,000 for steering.
@@
Figure 3: A fragment of a Python steering pair. The original code is 70 lines of text. The dict is the expected prediction. But, renaming config to __tmp0 makes the model mispredict Repository, which is a hallucination.
@@
For a given model $M$, our steering dataset is a set of triples $(x^{+},x^{-},t)$ where $x^{+},x^{-}$ are FIM type prediction prompts and $t$ is the expected type for both prompts. However, the maximum likelihood generation for $x^{+}$ is $M(x^{+})=t$ and for $x^{-}$ it is $M(x^{-})=t^{\prime}$, where $t^{\prime}\neq t$.
@@
To construct the negative prompt, we incrementally apply semantics-preserving code edits to the positive prompt until $M$ mispredicts $t$. We implement the following edits. 1) Rename variable:We select a variable binding (from a function definition) and rename it to an arbitrary name that does not conflict with other variables. We also rename all bound occurrences of the same variable so that the program’s semantics remain unchanged. 2) Remove type annotation:We select a type annotation (excluding the target $t$) and delete it. In a gradually typed language, removing or relaxing an annotation does not alter program semantics. 3) Rename user-defined type:We select an arbitrary type definition (e.g., a class name or a type alias) and rename it to an arbitrary name that does not conflict with other names in the program. 4) Rename builtin type:We also support renaming builtin types by introducing type aliases. Figure 2 illustrates several separate edits to a program.
By incrementally applying edits, the model eventually mispredicts $t$ (or else we fail to find an $x^{-}$ and thus discard $x^{+}$). Figure 3 illustrates a real example from our dataset where StarCoderBase-1B mispredicts after an edit. Note that a single edit often alters the positive prompt $x^{+}$ at several points. Moreover, when we apply several edits, the two prompts $x^{+},x^{-}$ may become significantly different from each other. We find that the mean edit-distance for our Python and TypeScript steering pairs is $74$ and $327$ respectively.
@@
Table 1: Overview of StarCoderBase-1B dataset sizes and applied code edits. Note that steering subsets are used for constructing steering vectors, while held out subsets are used to evaluate steering performance on out of distribution data. Dataset sizes vary because we only apply edits to programs for which the model predicts correct types. We then filter to remove overlap in source programs between held-out and steering sets.
Class balance
@@
We construct several datasets $\mathcal{D}$, where each dataset applies a subset (or all) of the aforementioned edits to programs in a language. We further split each dataset into a steering subset and a held-out evaluation subset. We ensure that the two subsets have prompts based on different source programs. This tests how well steering generalizes to wholly unseen programs. Table 1 summarizes the dataset sizes for StarCoderBase-1B. We report the sizes for StarCoderBase-7B in appendix A.
@@ 3.2 Constructing Steering Vectors
Given a dataset of steering pairs $(x_{i}^{+},x_{i}^{-},t)\in\mathcal{D}$ and model $M$, we apply forward passes $M(x_{i}^{+}),M(x_{i}^{-})$ and save model activations on the last token. Specifically, we extract activations from the residual stream of the model, which is the output of a transformer layer $\ell_{i}$ that in turn becomes the input to subsequent layers $\ell_{i+1...n}$. We write $A_{\ell}(x)$ to denote this activation tensor at some layer $\ell$ for prompt $x$. Thus we compute steering vectors $\mathbf{t}_{\ell}$—one for each layer—as the mean difference between positive and negative activations at that layer:
@@
The intuition behind equation 1 is that the distance between positive and negative pairs in activation space encodes the transformation for steering towards the correct type prediction.
@@
In the next section, we perform steering at the level of both individual layers and sets of adjacent layers simultaneously.
4 Results
In this section, we evaluate model performance on type prediction after activation steering. Note that performance before steering is 0.
4.1 At What Layers is Steering Most Effective?